### v1

@@ Abstract
Purpose: Dynamic positron emission tomography (PET) and
kinetic modeling are pivotal in advancing tracer development research in small
animal studies. Accurate kinetic modeling requires precise input function
estimation, traditionally achieved through arterial blood sampling. However,
arterial cannulation in small animals, such as mice, involves intricate,
time-consuming, and terminal procedures, precluding longitudinal studies. This
work proposes a non-invasive, fully convolutional deep learning-based approach
(FC-DLIF) to predict input functions directly from PET imaging data, which may
eliminate the need for arterial blood sampling in the context of dynamic
small-animal PET imaging.
Methods: The proposed FC-DLIF model consists of a spatial feature
extractor that acts on the volumetric time frames of the dynamic PET imaging
sequence, extracting spatial features. These are subsequently further processed
in a temporal feature extractor that predicts the arterial input function. The
proposed approach is trained and evaluated using images and arterial blood
curves from data using cross validation. Further, the model applicability
is evaluated on imaging data and arterial blood curves collected using two
additional radiotracers (, and ). The model was further evaluated
on data truncated and shifted in time, to simulate shorter, and shifted, PET
scans.
Results: The proposed FC-DLIF model reliably predicts the arterial
input function with respect to mean squared error and correlation. Furthermore,
the FC-DLIF model is able to predict the arterial input function even from
truncated and shifted samples. The model fails to predict the AIF from samples
collected using different radiotracers, as these are not represented in the
training data.
Conclusion: Our deep learning-based input function offers a non-invasive and reliable alternative to arterial blood sampling, proving robust and flexible to temporal shifts and different scan durations.
keywords:
dynamic positron emission tomography (PET), small-animal PET, arterial input function prediction, deep learning, prediction model1 Introduction
@@ 2 Materials and methods
This section describes the data, the design of the proposed deep learning model, and its training procedure. We refer the interested reader to the supplementary materials in Appendix A for more information on the data collection.
@@ 2.4 Evaluation metrics
The predicted DLIF curves were compared time frame by time frame with the respective measured AIF using a paired t-test (), and orthogonal regression to account for measurement errors in both predicted and measured variables. Normality was assessed using quantile-quantile plots. The coefficient of determination and Pearson correlation were also reported, describing the spread of the predicted curves. In order to assess the efficacy of the proposed method for downstream tasks, such as kinetic modeling, the graphical Patlak plot [30, 31] is used to produce net influx rates Ki given the reference and predicted input functions for each voxel over time. DLIF-based influx estimates were further compared to those obtained from the measured AIF using the same evaluation methods as for the predicted curves.
Comparisons were done against the DLIF model proposed by Kuttner et al. [22], which will be referred to as the baseline in the following sections. To ensure fair comparisons, the baseline was trained on the same data as the FC-DLIF, also using a -fold cross validation setting, but without data augmentation during training.
@@ 3.1 Quantitative results
Fig. 2 summarizes the distribution of mean squared error (MSE, Fig. 2(a)) and mean bias error (MBE, Fig. 2(b)) across all samples. FC-DLIF consistently achieved lower MSE values than the baseline, with a noticeably narrower spread and a more favorable distribution across percentiles. In terms of bias, FC-DLIF also reduced systematic under- or overestimation compared to the baseline, as reflected in the more symmetric MBE distribution centered around zero.
@@
Predictions from the baseline model are generally overestimating small values (less than ), and underestimating larger values, associated with the peak of the input function curve. The same charactheristics are not seen for the FC-DLIF method, but instead the spread is generally larger above values of , which is also shown in Fig. 9, for the peak labeled segments.
Downstream implications of AIF estimation accuracy are reflected in kinetic modeling results shown in Figure 4, where voxel-wise values derived from each model are compared to reference values obtained using the measured AIF and the Patlak model [33, 34]. FC-DLIF again yields higher agreement with the reference, confirming that improvements in input function estimation translate to better physiological parameter estimation.
Results on the unseen tracer data ( and ) showed decreased performance for both models, as expected due to tracer-specific uptake characteristics. Additional scatterplots and analysis are provided in the supplementary material (Figures 10 and 11).
@@ 4.1 Quantitative analysis
Although the precise source of performance gain is difficult to isolate, the two-step network inherently enforces a type of implicit regularization. This separation of concerns allows each subnetwork to specialize; spatial encoding in the SFE; and temporal modeling in the TFE, potentially introducing a form of structural regularization. This modular architecture may help the network learn more generalized temporal patterns in tracer dynamics, leading to improved alignment with the true AIF. The benefits are clearly reflected in the lower MSE and MBE distributions (Fig. 2) and the tighter correspondence between predicted and true AIF values (Fig. 3).
@@
The improvements in AIF prediction also translate into better physiological parameter estimation, as evidenced by the voxel-wise comparisons (Fig. 4). Accurate AIF estimation is critical for computing reliable kinetic parameters using models such as Patlak [33, 34], and FC-DLIF’s enhanced performance further supports its potential to replace invasive blood sampling without compromising modeling accuracy.
@@
When applied to the external dataset using tracers not seen during training ( and ), both models exhibited reduced performance. This result is consistent with the expectation that different tracers produce distinct uptake patterns and temporal kinetics, making generalization difficult without exposure during training. Future work may explore strategies for improving generalization across tracers, such as domain adaptation [35, 36], tracer-aware conditioning, or semi-supervised strategies [37].
@@ A.1 Animal experiments
The imaging data of mice used to train and evaluate the DLIF model were acquired at UiT The Arctic University of Norway (UiT). All animal experiments were approved by the Norwegian Food Safety Authority; FOTS id 29689. In total, healthy female mice from three different strains were included in this study: BALB/cJRj (N = ), C57BL/6JRj (N = ) and Balb/cAnNCrl (N = ), where BALB/cJRj made up the main group of samples. The remaining were set aside for evaluating the effects of the tracers (N = ) and (N = ). The mice were weeks of age upon arrival to the UiT animal facility, and fed ad libitum a standard rodent diet.
@@ A.2 Arterial and venous cannulation during image acquisition
At the time of PET/CT imaging, the mouse age was , with a corresponding weight of . The study used threedifferent radiotracers: , , and , where the mice receiving the radiotracer ( mice) were fasted for prior to injection; while the remaining—given and —were not fasted. The mice were administered anesthesia prior to the scan, weighed, and placed on a heated plate (), while receiving oxygen through a mask. A venous catheter was inserted into the tail vein of the mouse, for later radiotracer injection. The blood glucose during venous cannulation was measured using a glucose meter to . An incision in the neck of the mice enabled surgical cannulation of the carotid artery, allowing blood to be routed through a radiation detector at a withdrawal rate of . This setup facilitated concurrent measurements of whole blood activity during the PET scan with a temporal resolution of . To enable continuous arterial line measurements, an arterial-venous shunt was established, forming a closed loop between arterial sampling and venous reinfusion. Blood flowed sequentially from the arterial sampling line, radiation detector, a peristaltic pump, and a Y-connector—enabling intravenous radiotracer injection—before being reinfused via the venous injector, thereby preventing excessive blood loss.
The PET/CT imaging was performed using a -minute listmode scan on a TriumphTM LabPET-8TM small animal PET/CT scanner (TriFoil Imaging Inc., Chatsworth, CA, USA), and a sensor monitoring the respiration rate. The mice were injected with a using an automated injection pump, started after scanning was initiated. CT imaging was performed in order to correct for attenuation and scattering, after PET scanning. Following the scans, when the mice were still in deep anesthesia, they were euthanized using cervical dislocation. To account for variations in the PET scanner sensitivity, daily calibrations using a phantom were performed.
@@ A.3 Calibration and AIF processing
A calibration factor for the continuous arterial line measurements was derived as the ratio of the average signal from the continuous line measurements and the manual blood samples collected during the same blood sampling at each time point, corrected for delay. Calibration factors outside three scaled median absolute deviations from the median factors were considered outliers and discarded [39]. An overall calibration factor was determined as the average factor from the included blood sample factors. The sought AIF for each mouse was obtained by scaling the continuous line measurement data by the average calibration factor.
@@ B.1 Detailed results for individual imaging phases
Fig. 9 visualizes the distribution of error between the predicted input function from the baseline model [22] and FC-DLIF compared with the AIF, as a function of time since the start of tracer injection. A predominantly negative distribution in the plot would indicate the model underestimating the AIF.
@@ B.2 Results on unseen tracers dataset
Fig. 10 compares the baseline [22] and the proposed FC-DLIF models SUV predictions with the measured AIF. The lines are computed as the orthogonal regression line fitted to the point distributions of each respective color. A perfect predictor would have SUV points distributed across the black, dashed, line.
Fig. 11 use the predicted input function from either model in a Patlak graphical analysis [33, 34] to derive the kinetic parameters .
Fig. 12 shows some examples of input functions predicted by the proposed FC-DLIF model. The mean curve and the standard deviation over the runs are depicted for the best (left), the median (middle), and the worst (right) sample according to the MSE.

### v2

@@ Abstract
Dynamic positron emission tomography (PET) and kinetic modeling are pivotal in advancing tracer development research in small animal studies. Accurate kinetic modeling requires precise input function estimation, traditionally achieved through arterial blood sampling. However, arterial cannulation in small animals, such as mice, involves intricate, time-consuming, and terminal procedures, precluding longitudinal studies. This work proposes a non-invasive, fully convolutional deep learning-based approach (FC-DLIF) to predict input functions directly from PET imaging data, which may eliminate the need for arterial blood sampling in the context of dynamic small-animal PET imaging.
The proposed FC-DLIF model consists of a spatial feature extractor that acts on the volumetric time frames of the dynamic PET imaging sequence, extracting spatial features. These are subsequently further processed in a temporal feature extractor that predicts the arterial input function. The proposed approach is trained and evaluated using images and arterial blood curves from data using cross validation. Further, the model applicability is evaluated on imaging data and arterial blood curves collected using two additional radiotracers (, and ). The model was further evaluated on data truncated and shifted in time, to simulate shorter, and shifted, PET scans.
The proposed FC-DLIF model reliably predicts the arterial input function with respect to mean squared error and correlation. Furthermore, the FC-DLIF model is able to predict the arterial input function even from truncated and shifted samples. The model fails to predict the AIF from samples collected using different radiotracers, as these are not represented in the training data.
Our deep learning-based input function offers a non-invasive and reliable alternative to arterial blood sampling, proving robust and flexible to temporal shifts and different scan durations.
1 Introduction
@@ 2 Materials and methods
This section describes the data, the design of the proposed deep learning model, and its training procedure. Further details on data collection are provided in Appendix A.
@@ 2.4 Evaluation metrics
The predicted DLIF curves were compared time frame by time frame with the respective measured AIF using a paired t-test (), and orthogonal regression to account for measurement errors in both predicted and measured variables. Normality was assessed using quantile-quantile plots, reporting the Pearson correlation to describe the spread of the predicted curves. In order to assess the efficacy of the proposed method for downstream tasks, such as kinetic modeling, the graphical Patlak plot [30, 31] is used to produce net influx rates Ki given the reference and predicted input functions for each voxel over time. Prior to kinetic modeling, both the AIF and DLIF curves were converted to plasma input functions following [32]. DLIF-based influx estimates were further compared to those obtained from the measured AIF using the same evaluation methods as for the predicted curves.
Comparisons were done against the DLIF model proposed by Kuttner et al. [22], which will be referred to as the baseline in the following sections. To ensure fair comparisons, the baseline was trained on the same data as the FC-DLIF, also using a -fold cross validation setting.
@@ 3.1 Quantitative results
Fig. 2 summarizes the distribution of mean squared error (MSE, Fig. 2(a)) and injected noise (Fig. 2(b)) across all samples. FC-DLIF consistently achieved lower MSE values than the baseline, with a noticeably narrower spread and a more favorable distribution across percentiles. In terms of noise, FC-DLIF also demonstrated reduced variability, being less sensitive to noise in the input data.
@@
Predictions from the baseline model are generally overestimating small values (less than ), and underestimating larger values, associated with the peak of the input function curve. The same characteristics are not seen for the FC-DLIF method, but instead the spread is generally larger above values of , which is also shown in Fig. 9, for the peak labeled segments.
Downstream implications of AIF estimation accuracy are reflected in kinetic modeling results shown in Fig 4, where voxel-wise values derived from each model are compared to reference values obtained using the measured AIF and the Patlak model [34, 35]. FC-DLIF again yields higher agreement with the reference, confirming that improvements in input function estimation translate to better physiological parameter estimation.
Results on the unseen tracer data ( and ) showed decreased performance for both models, as expected due to tracer-specific uptake characteristics. Additional scatterplots and analysis are provided in the Appendix B (Figs. 10 and 11).
@@ 4.1 Quantitative analysis
Although the precise source of performance gain is difficult to isolate, the two-step network inherently enforces a type of implicit regularization. This separation of concerns allows each subnetwork to specialize; spatial encoding in the SFE; and temporal modeling in the TFE, potentially introducing a form of structural regularization. This modular architecture may help the network learn more generalized temporal patterns in tracer dynamics, leading to improved alignment with the true AIF. The benefits are clearly reflected in the lower MSE distribution (Fig. 2(a)), being less affected by a noisy input signal (Fig. 2(b)), and a tighter correspondence between predicted and true AIF values (Fig. 3).
@@
The improvements in AIF prediction also translate into better physiological parameter estimation, as evidenced by the voxel-wise comparisons (Fig. 4), and region-wise comparisons (Fig. 13(e) and Fig. 14(e)). Accurate AIF estimation is critical for computing reliable kinetic parameters using models such as Patlak [34, 35] and compartment models [36], and FC-DLIF’s enhanced performance further supports its potential to replace invasive blood sampling without compromising modeling accuracy.
@@
When applied to a dataset using tracers not seen during training ( and tracers), both models exhibited reduced performance. This result is consistent with the expectation that different tracers produce distinct uptake patterns and temporal kinetics, making generalization difficult without exposure during training. Future work may explore strategies for improving generalization across tracers, such as domain adaptation [37, 38], tracer-aware conditioning, or semi-supervised strategies [39].
@@ 5 Conclusion
Acknowledgments
This work is funded through Visual Intelligence, a Center for Research-based Innovation funded by the Research Council of Norway, grant no. 309439 and 303514. It was further supported by Tromsø Research Foundation (19_PET-NUKL) through the 180∘N Norwegian Nuclear Medicine Consortium and UiT Innovation funds (UiT Talent).
@@ A.1 Animal experiments
The imaging data of mice used to train and evaluate the FC-DLIF model were acquired at UiT The Arctic University of Norway (UiT). All animal experiments were approved by the Norwegian Food Safety Authority; FOTS id 29689. In total, healthy female mice from three different strains were included in this study: BALB/cJRj (N = ), C57BL/6JRj (N = ) and Balb/cAnNCrl (N = ), where BALB/cJRj made up the main group of samples. The remaining were set aside for evaluating the effects of the tracers (N = ) and (N = ). The mice were weeks of age upon arrival to the UiT animal facility, and fed ad libitum a standard rodent diet.
@@ A.2 Arterial and venous cannulation during image acquisition
At the time of PET/CT imaging, the mouse age was , with a corresponding weight of . The study used three different radiotracers: , , and . The mice receiving ( animals) were fasted for prior to injection, whereas the remaining animals—given and —were not fasted. The mice were anesthetized prior to the scan, weighed, and placed on a heated plate () while receiving oxygen through a mask. A venous catheter was inserted into the tail vein of each mouse for radiotracer injection. The blood glucose during venous cannulation was measured to . An incision in the neck enabled surgical cannulation of the carotid artery, allowing blood to be routed through a radiation detector at a withdrawal rate of . This setup facilitated concurrent measurements of whole blood activity during the PET scan with a temporal resolution of . To enable continuous arterial line measurements, an arterial–venous shunt was established, forming a closed loop between arterial sampling and venous reinjection. Blood flowed sequentially from the arterial sampling line, radiation detector, a peristaltic pump, and a Y-connector—enabling intravenous radiotracer injection—before being reinfused via the venous injector, thereby preventing excessive blood loss.
[…]
@@ A.3 Calibration and AIF processing
A calibration factor for the continuous arterial line measurements was derived as the ratio of the average signal from the continuous line measurements and the manual blood samples collected during the same blood sampling at each time point, corrected for delay. Calibration factors outside three scaled median absolute deviations from the median factors were considered outliers and discarded [41]. An overall calibration factor was determined as the average factor from the included blood sample factors. The arterial input function (AIF) for each mouse was obtained by scaling the continuous line measurement data by this average calibration factor.
@@ B.1 Detailed results for individual imaging phases
Figure 9 visualizes the distribution of error between the predicted input function from the baseline model [22] and FC-DLIF compared with the AIF, as a function of time since the start of tracer injection. A predominantly negative distribution in the plot would indicate the model underestimating the AIF.
@@ B.2 Results on unseen tracers dataset
Figure 10 compares the baseline [22] and the proposed FC-DLIF models’ SUV predictions with the measured AIF. The lines are computed as the orthogonal regression line fitted to the point distributions of each respective color. A perfect predictor would have SUV points distributed across the black, dashed, line.
Figure 11 uses the predicted input function from either model in a Patlak graphical analysis [34, 35] to derive the kinetic parameters .
Figure 12 shows examples of input functions predicted by the proposed FC-DLIF model. The mean curve and the standard deviation over the runs are depicted for the best (left), median (middle), and worst (right) sample according to the mean squared error (MSE).
B.3 Detailed kinetic modeling results
Further kinetic modeling results using the predicted input functions from both models are summarized in Figures 13 and 14. The kinetic parameters were estimated using an irreversible two-tissue compartment model [36] in two different regions: myocardium and brain. The estimated parameters were compared with those obtained using the measured AIF and DLIF. The quantile–quantile plots show the distribution of the kinetic parameters estimated using orthogonal regression.
Compared to the Patlak analysis presented in the main manuscript, the kinetic parameters derived from the two-tissue compartment model are less stable because the full measurement period is used in the fitting, and not only the linear phase. Outliers are therefore removed if they lie more than three standard deviations away from the mean of each parameter distribution. Once an outlier is detected, it is removed from all parameters in the respective region for consistency. This results in and samples remaining for the myocardium and brain regions, respectively.
The observed myocardium kinetic parameters (Figures 13 and 13(e)) follow the reference distribution well, with larger spread in , , and , as well as outliers not removed by the three standard deviation rule in and the blood volume fraction. Influx rates (Figure 13(e)) were higher than previously reported [42, 43]. This is likely due to shorter fasting times in our study, where longer fasting times are known to reduce influx rates in the myocardium [42, 43].
The parameter distributions in the brain (Figure 14) had fewer outliers, and similar to the myocardium, , , and had larger spread than the uptake and influx rate. The ranges of the brain kinetic parameters are in line with previous literature [44].